How long will it take a 2150 W motor to lift a 310 kg piano to a six story window locate 16.0 m a above ground? Express your answer in seconds.
Answer #1

Solution:

Given:

Power (P) = 2150 W

mass (m) = 310 kg

height (h) = 16 m

So, Work done (W) = m g h = (310 kg)(9.81 m/s2)(16 m) = 48657.6 J

Since : P = W / t

Thus: Time taken (t) = W / P = (48657.6 J) / (2150 W) = 22.63 sec